OpenStudy (anonymous):

Given that ABCDEFGH is a regular octagon and CDJK is a square find X which is the measure of angle CBK. Drawing is below.

OpenStudy (anonymous):



OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

In an octagon, the sum of the measures of the angles is (n - 2)180 = 6(180) = 1080 Each angle measures 1080/8 = 135 The octagon's side is congruent to all sides of the square, so BC = CK = y That means triangle BCK is isosceles, and angles CBK and CKB are congruent. Also, angle DCK is right angle and m<BCD = 135, so m<BCK = 135

OpenStudy (anonymous):

Okay, so would angle x = 22.5?

OpenStudy (mathstudent55):

right

OpenStudy (anonymous):

Thank you so much!

OpenStudy (mathstudent55):

wlcm